OpenStudy (anonymous):

Why do we multiply the 4 and 6? (a^4)^6

OpenStudy (anonymous):

Multiplication is just FAST ADDITION 3 x 4 = 3 + 3 + 3 + 3 = 12 but it is just easier to multiply 3 x 4 = 12 (a^4)^6 = (a^4) x (a^4) x (a^4) x (a^4) x (a^4) x (a^4) Now you could add all those 4's but it would be easier to multiply (a^4)^6 = a^24

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

law of exponent: power of a power (x^m)^n = x^mn (x^2)^3 = x^2×^3 = x^6 \[(a^{4})^{6}\] = (a*a*a*a)(a*a*a*a)(a*a*a*a)(a*a*a*a)(a*a*a*a)(a*a*a*a) \[=a ^{24}\]